# A probabilistic foundation model for crystal structure denoising, phase classification, and order parameters

### v1

@@ Introduction
Practically, we instantiate this idea by reusing the MACE-MP foundation model [Batatia2022mace, Batatia2023foundation] via transfer learning with fixed featurization on a curated subset of Materials Project structures [jain2013commentary] mapped to AFLOW prototypes [hicks2021aflow]. The training dataset includes elemental, binary, and ternary crystals, augmented with random elastic strains and Gaussian positional noise to mimic realistic thermal and mechanical perturbations. The resulting log-probability foundation model achieves near-perfect prototype classification and sub-Å denoising errors across hundreds of crystal types, maintains high accuracy on benchmark datasets of thermally perturbed configurations, and generalizes to challenging out-of-distribution scenarios such as shock-compressed Ti with coexisting BCC, HCP, FCC, and $\omega$ phases, as well as water–ice interfaces with mixed solid–liquid regions. These results demonstrate that a log-probability foundation model can provide a general, data-efficient route to automated structure recognition and probabilistic OPs for noisy atomistic configurations, while meeting the universality, robustness, and probabilistic interpretability criteria outlined above.
@@ Results
Figure 3 focuses on a silver structure in the hexagonal A_hP2_194 (space group 194) prototype. Panel 3a illustrates the effect of coupling or decoupling denoising from classification for a strongly perturbed configuration with Gaussian positional noise of 0.4 Å. When the model is applied in a purely classification mode (no denoising), the heavily distorted local environments cause the structure to be misclassified into a variety of competing prototypes. When denoising and classification are performed jointly through the shared $\log\!P$ landscape, the atomic positions are iteratively refined toward high-$\log\!P$ regions, and the model converges back to the correct A_hP2_194 class even from this strongly perturbed starting point.
@@
We next evaluate the robustness of the foundation model under realistic thermal disorder and local defects, where traditional geometric descriptors often struggle. For thermal effects, we use the DC3 database [chung2022data], which contains molecular dynamics (MD) snapshots of elemental and binary crystals equilibrated near their respective melting points. These configurations exhibit substantial vibrational amplitudes and partial loss of local order, providing a stringent test of structural recognition under near- and above-melting conditions.
@@
Figure 5 compares the classification performance of the log-probability foundation model against two widely used template-based methods, PTM and CNA. For each DC3 system, we take the highest-temperature snapshot available and consider a sequence of configurations obtained by applying $k=0,\dots,8$ denoising steps with the foundation model (with $k=0$ corresponding to the original DC3 snapshot). At each step $k$, we evaluate the log-probability model, PTM, and CNA on the same coordinates, i.e. on the configuration after $k$ denoising steps. Across all tested systems, our model reaches higher accuracy with fewer denoising iterations than PTM or CNA, and in many cases achieves 100% accuracy by step 3. At intermediate steps, where the structures remain visibly noisy, the log-probability model already outperforms both baselines, indicating that accurate structural recognition emerges early in the denoising process. This behavior reflects a key difference in philosophy: PTM and CNA rely on fixed geometric thresholds tuned to specific lattices, whereas the foundation model learns to associate a broad distribution of local environments with each prototype through its probabilistic training.
Our foundation model is not always the most accurate. For example, in the HCP panels of Fig. 5, the few-step classification accuracy falls behind PTM. This is because the model also predicts logits $l$ for many closely related close-packed prototypes such as A_hP3_191 and A_cF4_225, and with zero or few denoising steps the significant thermal perturbations lead to misclassifications into those competing phases. This is the price to pay for a foundation model that encompasses a large set of competing structures. In contrast, PTM considers only two close-packed phases, FCC and HCP. More broadly, it is conceivable that for specific applications involving only a few candidate phases, our foundation model may be less accurate than specialized classical or ML classifiers at zero or very few denoising steps. In such cases, one may either perform more denoising iterations or fine-tune the model on the restricted phase set.
Figure 5: Classification accuracy on thermally perturbed structures from the DC3 database. Top panels: accuracy versus denoising step for representative elemental and binary systems at high temperatures above their melting points, comparing the log-probability foundation model to PTM and CNA. The foundation model reaches higher accuracy within fewer denoising iterations and often achieves 100% accuracy by step 3. Bottom panel: summary of classification accuracy at step 2 across all tested systems, showing consistently superior early-stage performance over PTM and CNA under strong thermal disorder. For each denoising step $k$, PTM and CNA are evaluated on the same coordinates as the log-probability model, i.e. on the configuration obtained after $k$ log-probability denoising steps (with $k=0$ corresponding to the original DC3 snapshot).
@@
To probe sensitivity to point defects, we introduce vacancy-type defects into a BCC Fe supercell by randomly removing a small fraction of atoms (5 and 10 vacancies out of 432 atoms, respectively; Figure 6a and b). These missing atoms distort the local environments around the defect cores and frequently cause PTM to misclassify neighboring atoms as FCC or label them as “unknown” (left panels), reflecting the fragility of hard, template-based labels under local coordination changes. In contrast, the log-probability foundation model correctly assigns all atoms to the BCC prototype for both vacancy concentrations (middle panels), preserving the global phase identity. At the same time, the continuous BCC logit-based OP provides a natural defect-sensitive measure of local order: when atoms are colored by their BCC logit value $l_{a,\mathrm{BCC}}$ (right panels), the undisturbed crystal interior appears uniformly bright (high $l_{a,\mathrm{BCC}}$), while shells surrounding the vacancies show localized depressions in $l_{a,\mathrm{BCC}}$ (darker purple), indicating reduced confidence and stronger local disorder. Thus, the model simultaneously maintains robust global phase recognition and yields a smooth, quantitative measure of local deviations from ideal BCC order that discrete template matching cannot provide.
Figure 6: Classification and log-probability OP in defective BCC Fe. Two BCC Fe supercells with randomly removed atoms are used to create vacancy-type point defects: (a) 5 vacancies and (b) 10 vacancies out of 432 atoms. In each case, the left panel shows PTM classification, where atoms near the vacancy cores are frequently misclassified as FCC or labeled as “unknown,” reflecting the sensitivity of hard, template-based labels to local distortions. The middle panels show the log-probability foundation model classification via $\operatorname*{arg\,max}_{c}l_{ac}$, in which essentially all atoms are correctly identified as BCC. The right panels color atoms by $l_{a,\text{BCC}}$, with lighter purple indicating higher $l_{a,\text{BCC}}$ (bulk-like BCC environments) and darker purple indicating lower $l_{a,\text{BCC}}$ (locally perturbed environments near defects). The bulk interior remains uniformly light, while vacancy neighborhoods appear as darker halos, demonstrating that the model simultaneously recovers the correct global BCC prototype and provides a continuous, defect-sensitive OP that highlights local deviations from ideal crystalline order.
@@
Figure 7 illustrates two representative examples. Panel 7a shows an AgO structure in the AB_mP8_14 prototype, starting from a perturbed configuration and followed through successive denoising steps. At early iterations (e.g., step 1), some atoms are transiently assigned to alternative prototype classes such as A2B3_oF40_43, AB4_cP40_205, or AB2_cP6_224, reflecting local environments that momentarily resemble competing motifs. As denoising proceeds, these inconsistencies vanish and the model converges to a self-consistent assignment in which all atoms are correctly classified as the target AB_mP8_14 prototype.
@@
Panel 7b considers 5 distinct ZnO polymorphs. For each prototype (rows), we track the evolution of per-atom logit distributions across denoising steps (columns). The light curves correspond to the initial perturbed structures, intermediate curves show partially denoised states (step 3 of 8), and the darkest curves represent the fully denoised outputs. In all cases, the logit distributions for the true prototype sharpen into a dominant, well-separated peak, while the competing classes remain suppressed. Together, these examples demonstrate that the log-probability foundation model is not restricted to a small set of canonical lattices, but readily generalizes to diverse, low-symmetry AFLOW prototypes in binary systems.
@@
To probe the behavior of the log-probability foundation model in heterogeneous environments with coexisting ordered and disordered regions, we apply it to a water–ice interface featuring solid–liquid coexistence (Figure 8). The model is trained on 7 ordered ice polymorphs (Ic, Ih, II, III, VI, VII, and sI); liquid water is therefore out-of-distribution and expected to appear as low in all ice-related logits. The interface configuration is obtained from an equilibrated water–ice molecular dynamics simulation at 300 K and 1 kbar, so that thermal fluctuations naturally introduce positional disorder throughout the system. This setting is challenging for traditional local OPs because strong structural gradients and finite-temperature fluctuations blur the distinction between crystalline and liquid-like environments, particularly near the interface.
@@
As an extreme out-of-distribution test, we apply the log-probability foundation model to a large-scale simulation of shock-compressed Ti that exhibits severe deformation and complex phase coexistence (Figure 9). This configuration is not included in training and contains a heterogeneous mixture of close-packed (HCP-like) and $\omega$-like regions under highly nonequilibrium conditions, with additional highly strained environments that some classifiers label as BCC-like. This makes it a stringent benchmark for generalized structural recognition. An ablation study in which we retrain the model without elastic-strain augmentation shows that strongly shocked HCP Ti is then systematically misclassified as the rhombohedral AFLOW prototype A_hR3_166 (space group 166), confirming that strain augmentation is essential to avoid spuriously interpreting elastic distortions as phase changes (Supplementary Fig. S.2).
Figure 9: Shock-compressed Ti as an out-of-distribution test. (a) Classification using the full log-probability foundation model that includes chemical species. (b) Spatial map of the maximum per-atom logit over all phases from the full model, with vivid regions indicating atoms that strongly match a single prototype (high $\max_{c}l_{ac}$) and whitish regions indicating atoms whose environments are difficult to assign confidently to any single phase (low $\max_{c}l_{ac}$), e.g. near phase boundaries or in strongly distorted regions. (c) Classification using the chemistry-agnostic foundation model that ignores species and relies purely on geometry. (d) PTM classification baseline applied directly to the original shock-compressed configurations without log-probability denoising. Because PTM does not include an explicit $\omega$ prototype, structurally $\omega$-like regions are mapped onto HCP, BCC, or “unknown”, and large parts of the microstructure are either misclassified or left unassigned. Both log-probability foundation models recover HCP-like and $\omega$-like domains and resolve coexisting regions under strong strain and disorder, while PTM fails to reliably distinguish the $\omega$ phase and attributes many $\omega$-like atoms to BCC.
@@
Panel 9b illustrates how the logit field provides a confidence-based indicator for the classification. For each atom, we compute the maximum logit over all phases, $\max_{c}l_{ac}$; vivid regions correspond to atoms that strongly match a single prototype (high $\max_{c}l_{ac}$), while whitish regions have low $\max_{c}l_{ac}$ and are difficult to assign confidently to any phase. These low-$\max_{c}l_{ac}$ regions cluster around phase boundaries and highly distorted zones, naturally highlighting where the microstructure is structurally ambiguous or far from any ideal prototype.
@@ Acknowledgment
Appendix B Effect of elastic-strain augmentation on prototype classification

### v2

@@ Introduction
Practically, we instantiate this idea by reusing the MACE-MP foundation model [Batatia2022mace, Batatia2023foundation] via transfer learning with fixed featurization on a curated subset of Materials Project structures [jain2013commentary] mapped to AFLOW prototypes [hicks2021aflow]. The training dataset includes elemental, binary, and ternary crystals, augmented with random elastic strains and Gaussian positional noise to mimic realistic thermal and mechanical perturbations. The resulting log-probability foundation model achieves near-perfect prototype classification and sub-Å denoising errors across hundreds of crystal types, maintains high accuracy on benchmark datasets of thermally perturbed configurations, and generalizes to challenging out-of-distribution (OOD) scenarios such as shock-compressed Ti with coexisting BCC, HCP, FCC, and $\omega$ phases, as well as water–ice interfaces with mixed solid–liquid regions. These results demonstrate that a log-probability foundation model can provide a general, data-efficient route to automated structure recognition and probabilistic OPs for noisy atomistic configurations, while meeting the universality, robustness, and probabilistic interpretability criteria outlined above.
@@ Results
Figure 3 focuses on an Ag structure in the hexagonal A_hP2_194 (space group 194) prototype and is intentionally designed as an OOD probe of the coupled denoising–classification inference. While the model is trained with Gaussian positional noise amplitudes drawn uniformly up to $\sigma_{\max}=0.15$ Å (Methods), here we evaluate substantially larger perturbations, including $\sigma=0.4$ Å, to assess whether the learned $\log\hat{P}_{\theta}(\bm{r})$ landscape still provides a meaningful restoring drive toward the prototype manifold.
Panel 3a illustrates the qualitative difference between classification-only inference and the coupled denoising+classification inference for this strongly perturbed input ($\sigma=0.4$ Å). In classification-only mode (no denoising), the distorted local environments yield weak separation among competing prototypes, so no single class is strongly favored; consequently, atoms are distributed across multiple AFLOW labels. This behavior reflects a low-confidence near-tie regime rather than a confident but incorrect decision: at step 0 the per-phase logits occupy similar ranges and exhibit substantial overlap (panel 3c, top row). In contrast, when denoising is enabled, atomic positions are iteratively updated using the conservative score field $\hat{\bm{s}}(\bm{r})=\partial_{\bm{r}}\log\hat{P}_{\theta}(\bm{r})$ (Eq. 2), which drives the configuration toward higher-$\log\hat{P}_{\theta}$ regions and yields a self-consistent recovery of the A_hP2_194 assignment.
@@
We next evaluate robustness under ealistic thermal disorder and local defects, where traditional template- and threshold-based structure identifiers often struggle. For thermal effects, we use the DC3 database [chung2022data], which contains molecular dynamics (MD) snapshots of elemental and binary crystals equilibrated at high temperatures near their melting points. These configurations exhibit large vibrational amplitudes and, importantly, can also contain non-thermal disorder such as vacancies, interstitials, and stacking faults. Such environments are challenging for hard local classifiers because the local neighbor topology is no longer well represented by an ideal lattice template. The overall trends mirror our previous tests using the non-conservative denoiser model at a lower temperature[Hsu2024nCM-denoiser].
@@
Supplementary figure S.2 compares classification performance of the log-probability foundation model against two widely used baselines, PTM and CNA. For each DC3 system, we take the highest-temperature snapshot available and apply $k=0,\dots,8$ denoising steps using the foundation model (with $k=0$ corresponding to the original DC3 snapshot). At each step $k$, we evaluate all three methods on the same coordinates, i.e. on the configuration obtained after $k$ denoising steps. Across most tested systems, the foundation model attains higher accuracy with fewer denoising iterations than PTM or CNA, and in many cases reaches perfect phase identification within a few steps even when the structures remain visibly noisy. This reflects a key difference in philosophy: PTM and CNA rely on discrete, hand-crafted neighbor and topology criteria tuned to ideal lattices, whereas the foundation model learns a probabilistic association between a broad distribution of thermally perturbed local environments and their corresponding prototypes.
As a representative example, Fig. 5a shows BCC Li at $1.20\,T_{m}$. On the raw snapshot ($k=0$), the foundation model identifies BCC more reliably than PTM/CNA. As denoising proceeds, all methods improve when evaluated on the same denoised coordinates, but PTM/CNA retain a small fraction of non-BCC labels even at late steps.
A natural question is why PTM/CNA do not always reach 100% agreement with the reference label even after $k=8$ denoising steps for some systems (e.g., BCC Li/Fe). In addition to BCC Li, there are a few other cases in Supplementary Fig. S.2 with inconsistent classifications even at step 8. The reason is the presence of defects, e.g. vacancies, interstitials and Frenkel pairs, in these high temperature structures above $T_{m}$. Log-probability denoising is designed to suppress the approximately Gaussian thermal component while preserving such physically meaningful defect cores; consequently, the local environments near defects can remain far from any ideal template and may be labeled as “Other/Unknown” (or occasionally as a nearby lattice type) depending on the thresholds of PTM/CNA (Supplementary Fig. S.2a).
Our foundation model is not always the most accurate at very early denoising steps. In particular, for close-packed systems the few-step HCP accuracy can trail PTM. This is a consequence of the broader hypothesis space of the foundation model: it predicts logits over many closely related close-packed AFLOW prototypes (differing by stacking variants and subtle long-range order), which can be nearly degenerate under strong thermal disorder at $k=0$ or $k=1$. PTM, by contrast, typically distinguishes only a small set of close-packed templates (most commonly FCC vs. HCP). In applications where only a few phases are physically relevant, this gap can be mitigated by running additional denoising steps or by restricting inference to a reduced candidate prototype set.
@@
To directly probe defect sensitivity, we introduce vacancy-type defects into a BCC Fe supercell by randomly removing a small fraction of atoms (5 and 10 vacancies out of 432 atoms, respectively; Figure 5b and c). These missing atoms distort the local environments around the defect cores and frequently cause PTM to misclassify neighboring atoms as FCC or label them as “unknown” (left panels), reflecting the fragility of hard, template-based labels under local coordination changes. In contrast, the log-probability foundation model correctly assigns all atoms to the BCC prototype for both vacancy concentrations (middle panels), preserving the global phase identity. At the same time, the continuous BCC logit-based OP provides a natural defect-sensitive measure of local order: when atoms are colored by their BCC logit value $l_{a,\mathrm{BCC}}$ (right panels), the undisturbed crystal interior appears uniformly bright (high $l_{a,\mathrm{BCC}}$), while shells surrounding the vacancies show localized depressions in $l_{a,\mathrm{BCC}}$ (darker purple), indicating reduced confidence and stronger local disorder. Thus, the model simultaneously maintains robust global phase recognition and yields a smooth, quantitative measure of local deviations from ideal BCC order that discrete template matching cannot provide.
Figure 5: Robustness to thermal disorder and point defects. (a) Classification of BCC Li at $1.20~T_{m}$. The log-probability foundation model achieves higher accuracy than PTM/CNA on the raw high-temperature snapshot. Applying log-probability denoising improves all methods when evaluated on the denoised coordinates, but PTM/CNA typically plateau below 100% because vacancy/interstitial defects and other non-thermal disorder are preserved. (b,c) Defective BCC Fe with 5 and 10 vacancies (out of 432 atoms). PTM misclassifies atoms near vacancy cores as FCC or “unknown,” while the foundation model assigns BCC via $\operatorname*{arg\,max}_{c}l_{ac}$. Coloring by $l_{a,\mathrm{BCC}}$ reveals defect neighborhoods as low-logit halos.
@@
Figure 6 illustrates two representative examples. Panel 6a shows an AgO structure in the AB_mP8_14 prototype, starting from a perturbed configuration and followed through successive denoising steps. At early iterations (e.g., step 1), some atoms are transiently assigned to alternative prototype classes such as A2B3_oF40_43, AB4_cP40_205, or AB2_cP6_224, reflecting local environments that momentarily resemble competing motifs. As denoising proceeds, these inconsistencies vanish and the model converges to a self-consistent assignment in which all atoms are correctly classified as the target AB_mP8_14 prototype.
@@
Panel 6b considers 5 distinct ZnO polymorphs. For each prototype (rows), we track the evolution of per-atom logit distributions across denoising steps (columns). The light curves correspond to the initial perturbed structures, intermediate curves show partially denoised states (step 3 of 8), and the darkest curves represent the fully denoised outputs. In all cases, the logit distributions for the true prototype sharpen into a dominant, well-separated peak, while the competing classes remain suppressed. Together, these examples demonstrate that the log-probability foundation model is not restricted to a small set of canonical lattices, but readily generalizes to diverse, low-symmetry AFLOW prototypes in binary systems.
@@
To probe the behavior of the log-probability foundation model in heterogeneous environments with coexisting ordered and disordered regions, we apply it to a water–ice interface featuring solid–liquid coexistence (Figure 7). The model is trained on 7 ordered ice polymorphs (Ic, Ih, II, III, VI, VII, and sI); liquid water is therefore OOD and expected to appear as low in all ice-related logits. The interface configuration is obtained from an equilibrated water–ice molecular dynamics simulation at 300 K and 1 kbar, so that thermal fluctuations naturally introduce positional disorder throughout the system. This setting is challenging for traditional local OPs because strong structural gradients and finite-temperature fluctuations blur the distinction between crystalline and liquid-like environments, particularly near the interface.
@@
As an extreme OOD test, we apply the log-probability foundation model to a large-scale simulation of shock-compressed Ti that exhibits severe deformation and complex phase coexistence (Figure 8). This configuration is not included in training and contains a heterogeneous mixture of close-packed (HCP-like) and $\omega$-like regions under highly nonequilibrium conditions, with additional highly strained environments that some classifiers label as BCC-like. This makes it a stringent benchmark for generalized structural recognition. An ablation study in which we retrain the model without elastic-strain augmentation shows that strongly shocked HCP Ti is then systematically misclassified as the rhombohedral AFLOW prototype A_hR3_166 (space group 166), confirming that strain augmentation is essential to avoid spuriously interpreting elastic distortions as phase changes (Supplementary Fig. S.3).
Figure 8: Shock-compressed Ti as an OOD test. (a) Classification using the full log-probability foundation model that includes chemical species. (b) Spatial map of the maximum per-atom logit over all phases from the full model, with vivid regions indicating atoms that strongly match a single prototype (high $\max_{c}l_{ac}$) and pale regions indicate low-confidence/strongly distorted environments (low $\max_{c}l_{ac}$), e.g. near phase boundaries. (c) Classification using the chemistry-agnostic (geometry-only) foundation model. (d) PTM applied directly to the original shock-compressed configurations without log-probability denoising. Because PTM does not include an explicit $\omega$ prototype, structurally $\omega$-like regions are mapped onto HCP, BCC, or “unknown”. Both log-probability foundation models recover HCP-like and $\omega$-like domains and resolve coexisting regions under strong strain and disorder, while PTM often labels $\omega$-like atoms as BCC. Frame 75 was shown in (a-d). (e) Side-view snapshots from the shock-compression trajectory (frames 10, 15, 20, and 35), highlighting only $\omega$-classified atoms illustrating $\omega$ nucleation and growth as the shock propagates.
@@
Panel 8b illustrates how the logit field provides a confidence-based indicator for the classification. For each atom, we compute the maximum logit over all phases, $\max_{c}l_{ac}$; vivid regions correspond to atoms that strongly match a single prototype (high $\max_{c}l_{ac}$), while whitish regions have low $\max_{c}l_{ac}$ and are difficult to assign confidently to any phase. These low-$\max_{c}l_{ac}$ regions cluster around phase boundaries and highly distorted zones, naturally highlighting where the microstructure is structurally ambiguous or far from any ideal prototype.
@@ Acknowledgment
Appendix B Additional results on thermally perturbed DC3 structures
The DC3 database provides high-temperature MD snapshots that combine strong vibrational disorder with occasional non-thermal defects (e.g., vacancies/interstitials), posing a stringent test for local, template-based structure identifiers. Supplementary Fig. S.2 reports classification accuracy as a function of denoising step for representative elemental and binary systems. For each system, we start from the highest-temperature snapshot available ($k=0$) and apply $k=1,\dots,8$ log-probability denoising steps. To enable a like-for-like comparison, PTM and CNA are evaluated on the same coordinates at each step $k$ (i.e., on the configuration produced after $k$ log-probability denoising steps), so differences reflect the classifiers rather than differences in denoising. Overall, the foundation model reaches high accuracy with fewer denoising iterations, while template-based methods can plateau when defect-containing local environments remain difficult to match to ideal templates.
Figure S.2: Classification accuracy on thermally perturbed structures from the DC3 database. Top panels: accuracy versus denoising step for representative elemental and binary systems at high temperatures above their melting points, comparing the log-probability foundation model to PTM and CNA. The foundation model reaches higher accuracy within fewer denoising iterations and often achieves 100% accuracy by step 3. Bottom panel: summary of classification accuracy at step 2 across all tested systems, showing consistently superior early-stage performance over PTM and CNA under strong thermal disorder. For each denoising step $k$, PTM and CNA are evaluated on the same coordinates as the log-probability model, i.e. on the configuration obtained after $k$ log-probability denoising steps (with $k=0$ corresponding to the original DC3 snapshot).
Appendix C Effect of elastic-strain augmentation on prototype classification